P 13-15 (similar to) E Question Help All it ts financed 37% with dnt 5% with pre erred stock, and 58% with common stock lts cost of de is 5 a ts pre arra stok pays an annual dividend o 2 52 and is priced a $27 has an ฟืquity bata o 1 2 Assuma the nsk- rata 1s 1 %, the marka nsk pramium 15 nc 7% and AllCitys tax rate is 35% what is its after tax WACC? Note: Aszurme that the firrm will always be able to utilize it ful interest tax shield The WACC is )% (Round to two decimal places)
Answer #1

Cost of preferred stock = Annual dividend / preferred stock price

Cost of preferred stock = 2.52 / 27

Cost of preferred stock = 0.093 or 9.33%

Cost of equity using CAPM = Risk free rate + beta ( market risk premium)

Cost of equity = 1.6% + 1.2 ( 6.7%)

Cost of equity = 9.64%

WACC = Weight of stock*cost of stock + weight of preferred stock*cost of preferred stock + weight of debt*after tax cost of debt

WACC = 0.58*0.0964 + 0.05*0.093 + 0.37*0.058*(1 - 0.35)

WACC = 0.055912 + 0.00465 + 0.013949

WACC = 0.0745 or 7.45%